OpenStudy (anonymous):

I don't understand, I've tried several times. An auditorium earned $30,000 in sold-out concert ticket sales. Front section tickets cost $75 per seat and back section tickets cost $50 per seat. The number of front section seats is twice the number of back section seats. How many seats are in the front section?

OpenStudy (anonymous):

\(\$50B+\$75F=\$30,000\) \(F=2B\)

OpenStudy (anonymous):

it would be 8 right?

OpenStudy (anonymous):

If \(F=2B\) then \(B=\frac{F}{2}\) \(\$50\left(\frac{F}{2}\right)+\$75F=\$30,000\) \(\$100 F=\$30,000\)